# references.bib

### v1

@@ III-A Types of Representations
Understanding 3D representation types is crucial for evaluating reconstruction methods in robotics, as each representation influences downstream applications differently. Mesh representations are particularly important for physics simulation in robotics, while point clouds are common from RGB-D sensors. Representation choice affects memory requirements, computational efficiency, and rendering quality -- all critical factors for on-robot deployment. Importantly, current robotics simulators are optimised for specific representations (primarily meshes), making representation compatibility a practical necessity rather than just a theoretical concern. 3D representation can be broadly split into two categories: explicit and implicit representations. As some 3D reconstruction models use intermediate representations of different modalities, we briefly explain examples of both explicit and implicit representations.
@@
Explicit representations store each data point individually. Point clouds, for example, define surfaces with sets of 3D coordinates $\{x|x\in\mathbb{R}^{3}\}$ in Euclidean space. Due to the proliferation of RGB-D cameras in robotics [tadic_perspectives_2022], point clouds are often readily available in robotics domains and even though they frequently produce artifacts such as holes, they provide comparably more accurate information about scene geometry than 2D RGB projections. Several works propose methods for point cloud estimation or completion [lin_learning_2018, mandikal_dense_2019, melas-kyriazi_pc2_2023, nichol_point-e_2022]. For a detailed overview of 3D reconstruction with point clouds, see [huang_surface_2024]. Meshes are among the most widely used 3D representations for simulation as they allow efficient rendering and accurate rigid body collision computation. As a result, simulators are primarily built around mesh representations [coumans_pybullet_2016, makoviychuk_isaac_2021, mittal_orbit_2023]. Meshes consist of sets of vertices $V=\{v|v\in\mathbb{R}^{3}\}$, and faces $F=\{(v_{n})_{n=0}^{N}|v_{n}\in V\}$, surfaces that span subsets of vertices and define the object surface. Due to this very specific structure, directly predicting meshes is non-trivial and existing methods mostly resort to predicting deformation fields on mesh initialisations (e.g. [pontes_image2mesh_2019, jack_learning_2019, wang_pixel2mesh_2018]), or predicting other representations that are subsequently converted into meshes through techniques like ray marching or marching cubes (e.g. [liu_zero-1--3_2023, tochilkin_triposr_2024, boss_sf3d_2024, liu_one-2-3-45_2023]). Exceptions include [siddiqui_meshgpt_2024], who define a sequential order on the mesh vertices and faces to predict mesh structures with a transformer architecture [vaswani_attention_2017]. Mesh reconstruction from visual observations has received significant attention from the computer vision community [wang_pixel2mesh_2018, groueix_papier-mache_2018, mezghanni_physically-aware_2021, liu_zero-1--3_2023, qian_magic123_2023, chen_fantasia3d_2023, liu_one-2-3-45_2023, boss_sf3d_2024, tang_lgm_2025, xu_instantmesh_2024] and is the main focus of our evaluations in this work due to the widespread support for meshes in robotic simulators. Gaussian Splatting [kerbl_3d_2023] is a relatively recent addition to explicit representations. Each Gaussian is defined by its centre position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ defining the orientation and shape of the Gaussian blob, an opacity value $\sigma\in[0,1]$, and its colour. Together, this set of Gaussian blobs defines object surfaces or volumes. Similar to other explicit representations, Gaussian Splatting representations are very efficient to render but come with memory overhead as each Gaussian needs to be stored individually. Due to its high visual fidelity, Gaussian Splatting has recently become a focus for 3D reconstruction [melas-kyriazi_im-3d_2024, szymanowicz_splatter_2024, yan_phycage_2024, xu_grm_2025, xu_agg_2024, szymanowicz_flash3d_2024, liu_gaussian_2024].
@@
Implicit representations instead define a function $F:\mathbb{R}^{3}\rightarrow I$, where 3D coordinates $x\in\mathbb{R}^{3}$ are mapped to elements in $I$. $I$ can have various definitions, such as occupancy $\sigma\in[0,1]$, colour $c\in[0-255]^{3}$, signed distance $s\in\mathbb{R}$, or any combination of these. The number of parameters in $F$ can be set strictly smaller than the number of coordinates that $F$ is queried with, resulting in a (potentially lossy) compression of the geometric and visual information. At the same time, however, rendering is very inefficient as $F$ has to be queried many times to accumulate geometry and visual information along the camera rays. Specific implicit representation models include Neural Radiance Fields [mildenhall_nerf_2022] and DeepSDF [park_deepsdf_2019], and Occupancy Networks [mescheder_occupancy_2019].
@@ III-B Input Data
3D reconstruction models can operate on a wide variety of input data modalities. We are concerned with 3D reconstruction in a typical robotics context and focus on reconstruction from image or depth data. For an overview of 3D reconstruction from other modalities such as text, see [li_generative_2023]. 3D reconstruction approaches can be categorised based on the density of input viewpoints required: dense-view, sparse-view, and single-view methods, each with distinct characteristics and applications in robotics scenarios.
@@
Dense-view 3D reconstruction, that is reconstructing only what is observed, has a rich history. If camera poses, projection parameters, and depth data are available, it is straightforward to convert depth information to dense point clouds in the world coordinate system. Without camera extrinsics, reconstruction becomes an optimisation problem. One major line of work, Simultaneous Localisation and Mapping (SLAM), concerns reconstruction of a map of the environment and estimation of the camera trajectory in it from real-time sensor data [aulinas_slam_2008]. Another line of work, Structure-from-Motion, does not operate on online sequential data but instead reconstructs 3D structures and estimates camera viewpoints from unordered sets of offline images [schonberger_structure--motion_2016]. 3D Gaussian Splatting [kerbl_3d_2023], a more recent form of dense-view reconstruction, usually relies on unordered sets of dense views with camera extrinsics available.
@@
Sparse-view reconstruction alleviates the need for complete scene coverage in the reconstruction inputs by leveraging explicit priors, exploiting artifacts in scene representations, or relying on learnt parametric models to complete unobserved regions. [zheng_beyond_2013] utilise explicit priors for volumetric completion while [li_rico_2023] makes assumptions on object smoothness and object-background relations. [kong_vmap_2023] exploits artifacts resulting from implicit neural representations: as neural networks tend towards smooth function approximation, they observe that implicit neural representations naturally perform object closure and completion to some degree. [yu_monosdf_2022] use parametric models to predict depth and surface normal maps for input images and report increased detail in reconstruction when using these cues. Other works learn parametric models that individually embed multiple views in latent space, fuse these embeddings, and decode towards complete 3D reconstructions [choy_3d-r2n2_2016, kar_learning_2017, xie_pix2vox_2019]. [liu_gaussian_2024] cluster 3D Gaussians into individual object and complete occluded surfaces using a variational autoencoder [kingma_auto-encoding_2014]. [charatan_pixelsplat_2024] learn a feed-forward model to reconstruct radiance fields from two views and [chen_single-stage_2023] learn a density function over NeRFs that allows image-conditioned sampling with arbitrary numbers of views.
@@
Single-view reconstruction is arguably more difficult for producing faithful reconstructions as substantial portions of a scene are unobserved. Due to this ill-position, proposed methods are trained on datasets to capture general geometric properties of the objects in our world. Single-view reconstruction is particularly central to our work, as it represents the most practical and generalisable scenario for robotic deployment. Unlike multi-view methods that require substantial scene coverage or specialised equipment, single-view approaches align with real-world constraints where robots must rapidly assess and interact with novel environments from limited viewpoints. While these methods traditionally face greater challenges in producing accurate reconstructions, their successful development would enable transformative capabilities for autonomous systems operating in unstructured environments. Broadly speaking, methods can be categorised into two distinct flavours. Firstly, and historically more prominently, there are methods that only rely on the single input view during 3D reconstruction by either directly regressing 3D shapes (for example [fan_point_2017, szymanowicz_splatter_2024, chen_transformers_2022, li_instant3d_2023]) or via test-time optimisation (for example [pavllo_shape_2023, lee_understanding_2022, jain_zero-shot_2022]).
@@
The second popular flavour of single-view 3D reconstruction is leveraging multi-view generative models in the reconstruction process. One way of doing so is explicitly generating additional views based on the input image using multi-view diffusion models, predicting 3D Gaussians for each individual view using a U-Net [tang_lgm_2025] or transformer [xu_grm_2025], and finally fusing the predicted Gaussians. [melas-kyriazi_im-3d_2024] instead use generated views as rendering targets in the original Gaussian Splatting formulation [kerbl_3d_2023]. [xu_instantmesh_2024] choose to predict triplane features from multiple artificial views and extract meshes while [liu_one-2-3-45_2023] use a dense set of artificially generated views to predict signed distance values. Multi-view consistency is generally an issue in these approaches as any inconsistencies directly translate to artifacts in the reconstruction. [liu_one-2-3-45_2023] address this by using intermediate cost-volumes that capture inconsistencies. [liu_one-2-3-45_2024] instead fine-tune a multi-view generative model to be more consistent across views, and [liu_syncdreamer_2023] and [shi_mvdream_2023] propose to condition on a joint feature space derived from multiple views. Some approaches utilise additional information such as albedo or surface normal maps to increase reconstruction fidelity [chen_2l3_2024, wu_unique3d_2024]. A separate line of work does not use explicit multi-view generation but instead uses multi-view diffusion models to guide optimisation of an underlying 3D representation [liu_zero-1--3_2023, qian_magic123_2023]. This is explained in more detail in the following Section III-C.
@@ III-C Optimisation Paradigms
3D reconstruction can leverage several approaches to arrive at their outputs. Given our interest in robotic applications, the choice of optimisation paradigm is particularly relevant as it directly impacts computational requirements, inference speed, and ultimately the feasibility of on-device deployment. Models with heavy computational demands may be impractical for real-time robotic applications with limited on-board computing resources. In this section, we highlight two primary approaches: feed-forward prediction and inference-time optimisation.
@@
Feed-forward prediction is most in line with the standard use-case of deep learning architectures. For example, [wang_pixel2mesh_2018] propose a series of convolutional layers to gradually deform ellipsoid mesh initialisations, [fan_point_2017] propose an encoder-decoder structure to predict point cloud coordinates. For a detailed overview of such reconstruction methods up to the year 2021, please refer to [han_image-based_2021]. More recently, other models use feed-forward prediction to regress 3D Gaussians: [szymanowicz_splatter_2024] use a U-Net [ronneberger_u-net_2015] to predict one Gaussian blob per input image pixel while allowing arbitrary position offsets for 3D Gaussians. The simplicity of this architecture enables very fast prediction at up to 38 Hz. [xu_agg_2024] instead use transformers to predict 3D Gaussians based on DINOv2 [oquab_dinov2_2023] image features and [szymanowicz_flash3d_2024] rely on a ResNet-type [he_deep_2016] backbone. Implicit representations can also be produced in a feed-forward manner: [chen_transformers_2022] learn to regress the parameters of NeRFs using an image-conditioned transformer network while [jun_shap-e_2023] train a diffusion model on NeRF MLP weight matrices. [hong_lrm_2023] propose a model that maps from image input to triplane features which, when interpolated, can be used to query radiance information at spatial locations. [tochilkin_triposr_2024] and [boss_sf3d_2024] propose several improvements including data curation, masked losses for higher reconstruction fidelity, material and illumination estimation, and efficient mesh extraction.
@@
A major limitation of feed-forward shape prediction is that 3D datasets have historically been orders of magnitude smaller than their 2D counterparts. This has often limited feed-forward reconstructions to few specific object categories [choy_3d-r2n2_2016, wang_pixel2mesh_2018, rematas_sharf_2021] as any attempt at open-vocabulary reconstruction would be too far removed from training distributions. In response to this, several works proposed leveraging pre-trained 2D vision models in a GAN-style [goodfellow_generative_2014] setup -- training a 3D generative model with guidance through 2D renderings [nguyen-phuoc_hologan_2019, zhang_image_2020, rajeswar_pix2scene_2018]. Simultaneously, significant efforts have been made towards scaling up 3D datasets [deitke_objaverse-xl_2023] which has lead to impressive feed-forward reconstruction performances [huang_zeroshape_2024, xu_dmv3d_2023, li_instant3d_2023, wang_pf-lrm_2023, zou_triplane_2024].
@@
Inference-time optimisation of an underlying 3D representation with guidance signals from pre-trained 2D vision models is an alternative to feed-forward prediction. Some works propose using CLIP [radford_learning_2021] to produce gradients on 2D renderings of the underlying 3D asset [hong_avatarclip_2022, khalid_clip-mesh_2022, lee_understanding_2022, jain_zero-shot_2022]. If the rendering process itself is differentiable, the chain-rule can be used to obtain gradient signals w.r.t. the 3D representation. However, the bag-of-words behaviour of CLIP models [yuksekgonul_when_2022] usually limits the visual fidelity. More recently, powerful 2D diffusion models have replaced CLIP in guiding shape optimisation. [poole_dreamfusion_2022] and [wang_score_2023] both propose a process that repeatedly adds noise to 2D renders of an underlying 3D asset, then uses a pre-trained 2D diffusion model to obtain the 2D score (the gradient of the log-density over the noised 2D image space w.r.t. the 2D image), and apply the chain rule from 2D image to 3D asset to update the asset. An issue with these early diffusion approaches is that the diffusion score cannot be sufficiently guided for multi-view usage -- a requirement for making meaningful updates to all parts of the 3D asset. In response, [liu_zero-1--3_2023] train a view-conditioned 2D diffusion model and increase the 3D reconstruction quality significantly. [qian_magic123_2023] propose several extensions, including using both 2D and 3D diffusion models and a multi-stage setup. Generally, runtime and memory consumption is a major concern with these approaches, with object generations often requiring multiple hours of optimisation.
III-D Representation Abstractions
@@
3D reconstruction approaches can be broadly categorised based on their representational abstraction level. Wholistic approaches, including many SLAM systems, typically focus on generating unified representations of entire scenes (e.g. [wang_neuris_2022, yu_monosdf_2022, szymanowicz_flash3d_2024, turkulainen_dn-splatter_2025, szymanowicz_bolt3d_2025]). These methods generally rely on depth information or estimation to reconstruct observed elements, resulting in representations that essentially project 2.5D information into 3D space.
@@
In contrast, object-centric approaches focus on reconstructing individual objects as distinct entities [wang_pixel2mesh_2018, melas-kyriazi_im-3d_2024, xu_instantmesh_2024, tang_lgm_2025, li_rico_2023]. These methods prioritise resolving self- and viewpoint-occlusion through heuristics or parametric models trained to capture typical object features from datasets.
@@
Between these two extremes lies an emerging middle ground: generating digital twins of entire scenes by fully reconstructing constituent objects while maintaining their individuality [yao_cast_2025, dogaru_generalizable_2024, huang_midi_2024, han_reparo_2024]. This hybrid approach is particularly valuable for robotics applications as it enables both scene-level understanding and object-level interaction. Additionally, it allows reconstruction algorithms to leverage contextual information about an object’s surroundings, providing valuable cues about potential occlusions and spatial relationships. For example, proximity between objects suggests possible unseen object parts due to occlusion, rather than assuming these parts simply do not exist because they are not visible in the observation.

### v2

@@ III-A Types of Representations
Explicit representations store each data point individually. Point clouds, for example, define surfaces with sets of 3D coordinates in Euclidean space. Due to the proliferation of RGB-D cameras [67], point clouds are often readily available in robotics domains and provide comparably more accurate information about scene geometry than 2D RGB projections.
@@
For a detailed overview of 3D reconstruction with point clouds, see [86]. Meshes are among the most widely used 3D representations for simulation as they allow efficient rendering and accurate rigid body collision computation. As a result, simulators are primarily built around mesh representations [41, 42, 87]. Meshes consist of sets of vertices and faces, surfaces that span subsets of vertices and define the object surface. Due to this very specific structure, directly predicting meshes is non-trivial and existing methods mostly resort to predicting deformation fields on mesh initialisations (e.g. [88, 89, 90]), or predicting other representations that are subsequently converted into meshes through techniques like ray marching or marching cubes (e.g. [70, 68, 80, 26]). Exceptions include [91], who define a sequential order on the mesh vertices and faces to predict mesh structures with a transformer architecture [92]. Mesh reconstruction from visual observations has received significant attention from the computer vision community [90, 93, 94, 70, 79, 95, 26, 80, 96, 27] and is the main focus of our evaluations in this work due to the widespread support for meshes in robotic simulators. 3DGS [43] is a relatively recent addition to explicit representations. Each Gaussian is defined by its centre position in Euclidean space, a covariance matrix defining orientation and shape of the Gaussian blob, an opacity value, and its colour. Together, a set of Gaussian blobs defines object surfaces or volumes. 3DGS representations are very efficient to render but come with memory overhead as each Gaussian needs to be stored individually. Due to its high visual fidelity, 3DGS has recently become a focus for novel view synthesis [43, 97] and 3D reconstruction [98, 99, 75, 100, 101, 102, 103, 104, 105].
@@
Implicit representations instead define a parametric function where 3D coordinates are mapped to various quantities of interest, such as occupancy, colour, or signed distance. The number of free parameters can be set strictly smaller than the number of Euclidean coordinates, resulting in a (potentially lossy) compression of the geometric and visual information. At the same time, however, rendering is very inefficient as the function has to be queried many times to accumulate geometry and visual information along the camera rays. Specific models include DeepSDF [106] and Occupancy Networks [107].
@@ III-B Input Data
We focus on reconstruction from image or depth data in robotics contexts. Such 3D reconstruction approaches can be categorised based on the density of input viewpoints required: dense-view, sparse-view, and single-view methods, each with distinct characteristics and applications in robotics scenarios.
@@
Dense-view 3D reconstruction, that is reconstructing only what is observed, has a rich history. If camera poses, projection parameters, and depth data are available, it is straightforward to convert depth information to dense point clouds in the world coordinate system. Without camera extrinsics, reconstruction becomes an optimisation problem. One major line of work, Simultaneous Localisation and Mapping (SLAM), concerns reconstruction of a map of the environment and estimation of the camera trajectory in it from real-time sensor data [108]. Another line of work, Structure-from-Motion, does not operate on online sequential data but instead reconstructs 3D structures and estimates camera viewpoints from unordered sets of offline images [109]. 3DGS [43], a more recent form of dense-view reconstruction, typically relies on unordered sets of dense views with camera extrinsics available; extensions to simultaneous localisation and mapping have also been proposed [110, 111, 112].
@@
Sparse-view reconstruction alleviates the need for complete scene coverage in the reconstruction inputs by leveraging explicit priors [113, 78], exploiting artifacts in scene representations [77], or relying on learnt parametric models [114, 115, 116, 117] to complete unobserved regions. [113] utilise explicit priors for volumetric completion while [78] makes assumptions on object smoothness and object-background relations. [77] exploits artifacts resulting from implicit neural representations: as neural networks tend towards smooth function approximation, they observe that implicit neural representations naturally perform object closure and completion to some degree. [114] use parametric models to predict depth and surface normal maps for input images and report increased detail in reconstruction when using these cues. Other works learn parametric models that individually embed multiple views in latent space, fuse these embeddings, and decode towards complete 3D reconstructions [115, 116, 117]. [103] cluster 3D Gaussians into individual object and complete occluded surfaces using a variational autoencoder [118]. [119] learn a feed-forward model to reconstruct radiance fields from two views and [120] learn a density function over NeRFs that allows image-conditioned sampling with arbitrary numbers of views.
@@
Single-view reconstruction is arguably more difficult for producing faithful reconstructions as substantial portions of a scene are unobserved. Due to this ill-position, proposed methods are trained on datasets to capture general geometric properties of the objects in our world. Single-view reconstruction is particularly central to our work, as it represents the most practical and generalisable scenario for robotic deployment. Unlike multi-view methods that require substantial scene coverage or specialised equipment, single-view approaches align with real-world constraints where robots must rapidly assess and interact with novel environments from limited viewpoints. While these methods traditionally face greater challenges in producing accurate reconstructions, their successful development would enable transformative capabilities for autonomous systems operating in unstructured environments. Broadly speaking, methods can be categorised into two distinct flavours. Firstly, and historically more prominently, there are methods that only rely on the single input view during 3D reconstruction by either directly regressing 3D shapes (for example [121, 99, 122, 123]) or via test-time optimisation (for example [124, 125, 126]).
@@
The second popular flavour of single-view 3D reconstruction is leveraging multi-view generative models in the reconstruction process. One way of doing so is explicitly generating additional views based on the input image using multi-view diffusion models, predicting 3D Gaussians for each individual view using a U-Net [96] or transformer [100], and finally fusing the predicted Gaussians. [98] instead use generated views as rendering targets in the original 3DGS formulation [43]. [27] choose to predict triplane features from multiple artificial views and extract meshes while [26] use a dense set of artificially generated views to predict signed distance values. Multi-view consistency is generally an issue in these approaches as any inconsistencies directly translate to artifacts in the reconstruction. [26] address this by using intermediate cost-volumes that capture inconsistencies. [81] instead fine-tune a multi-view generative model to be more consistent across views, and [69] and [127] propose to condition on a joint feature space derived from multiple views. Some approaches utilise additional information such as albedo or surface normal maps to increase reconstruction fidelity [128, 129]. A separate line of work does not use explicit multi-view generation but instead uses multi-view diffusion models to guide optimisation of an underlying 3D representation [70, 79]. This is explained in more detail in the following Section III-C.
@@ III-C Optimisation Paradigms
In this section, we highlight two primary approaches 3D reconstruction leverage to arrive at their outputs: feed-forward prediction and inference-time optimisation.
@@
Feed-forward prediction is most in line with the standard use-case of deep learning architectures. For example, [90] propose a series of convolutional layers to gradually deform ellipsoid mesh initialisations, [121] propose an encoder-decoder structure to predict point cloud coordinates. For a detailed overview of such reconstruction methods up to the year 2021, please refer to [130]. More recently, other models use feed-forward prediction to regress 3D Gaussians: [99] use a U-Net [131] to predict one Gaussian blob per input image pixel while allowing arbitrary position offsets for 3D Gaussians. The simplicity of this architecture enables very fast prediction at up to 38 Hz. [101] instead use transformers to predict 3D Gaussians based on DINOv2 [132] image features and [102] rely on a ResNet-type [133] backbone. Implicit representations can also be produced in a feed-forward manner: [122] learn to regress the parameters of NeRFs using an image-conditioned transformer network while [134] train a diffusion model on NeRF MLP weight matrices. [135] propose a model that maps from image input to triplane features which, when interpolated, can be used to query radiance information at spatial locations. [68] and [80] propose several improvements including data curation, masked losses for higher reconstruction fidelity, material and illumination estimation, and efficient mesh extraction.
@@
A major limitation of feed-forward shape prediction is that 3D datasets have historically been orders of magnitude smaller than their 2D counterparts. This has often limited feed-forward reconstructions to few specific object categories [115, 90, 136], since the limited data volume precludes effective training for category-agnostic reconstruction. In response to this, several works proposed leveraging pre-trained 2D vision models in a GAN-style [137] setup - training a 3D generative model with guidance through 2D renderings [138, 139, 140]. Simultaneously, significant efforts have been made towards scaling up 3D datasets [141, 48] which has lead to impressive feed-forward reconstruction performances [142, 143, 123, 144, 145].
@@
Inference-time optimisation of an underlying 3D representation with guidance signals from pre-trained 2D vision models is an alternative to feed-forward prediction. Some works propose using CLIP [146] to produce gradients on 2D renderings of the underlying 3D asset [147, 148, 125, 126]. If the rendering process itself is differentiable, the chain-rule can be used to obtain gradient signals w.r.t. the 3D representation. However, the bag-of-words behaviour of CLIP models [149] usually limits the visual fidelity. More recently, powerful 2D diffusion models have replaced CLIP in guiding shape optimisation. [50] and [49] both propose a process that repeatedly adds noise to 2D renders of an underlying 3D asset, then uses a pre-trained 2D diffusion model to obtain the 2D score (the gradient of the log-density over the noised 2D image space w.r.t. the 2D image), and apply the chain rule from 2D image to 3D asset to update the asset. An issue with these early diffusion approaches is that the diffusion score cannot be sufficiently guided for multi-view usage - a requirement for making meaningful updates to all parts of the 3D asset. In response, [70] train a view-conditioned 2D diffusion model and increase the 3D reconstruction quality significantly. [79] propose several extensions, including using both 2D and 3D diffusion models and a multi-stage setup. Generally, runtime and memory consumption is a major concern with these approaches, with object generations often requiring multiple hours of optimisation.
III-D Scene-level vs. Object-level Reconstruction
@@
3D reconstruction approaches can be broadly categorised based on their representational abstraction level. Wholistic approaches, including many SLAM systems, typically focus on generating unified representations of entire scenes (e.g. [150, 114, 102, 151, 152]). These methods generally rely on depth information or estimation to reconstruct observed elements, resulting in representations that essentially project 2.5D information into 3D space.
@@
In contrast, object-centric approaches focus on reconstructing individual objects as distinct entities [90, 98, 27, 96, 78]. These methods prioritise resolving self- and viewpoint-occlusion through heuristics or parametric models trained to capture typical object features from datasets.
@@
Between these two extremes lies an emerging middle ground: generating digital twins of entire scenes by fully reconstructing constituent objects while maintaining their individuality [36, 153, 154, 155]. This hybrid approach is particularly valuable for robotics applications as it enables both scene-level understanding and object-level interaction. Additionally, it allows reconstruction algorithms to leverage contextual information about an object’s surroundings, providing valuable cues about potential occlusions and spatial relationships. For example, proximity between objects suggests possible unseen object parts due to occlusion, rather than assuming these parts simply do not exist because they are not visible in the observation.